OpenStudy (anonymous):

A cylinder of mass 65kg has a radius of gyration 0.8m . the cylinder is rotated by a belt of at 4rev/s. calculate the rotational kinetic energy of the cylinder

OpenStudy (anonymous):

rotational KE is given by kE=1/2Iw^2 w = 4rev/s = 8(pi)/s I = m(Rg)^2 = 65*.08^2 If the I is right then you should be able to cont from here.. if not, sorry :)